Suppose a firm that produces light bulbs wants to know whether it can say that its light bulbs typically last more than 1500 hours. Hoping to find support for their claim, the firm collects a random sample of n = 25 light bulbs and records the lifetime (in hours) of each bulb. The information related to the hypothesis test is presented below.
Test of H0: 1500 versus H1: > 1500
Sample mean 1509.5
Std error of mean 4.854
What is the test value that you would use to conduct this test

Found 2 solutions by ewatrrr, ed42ptt:
Answer by ewatrrr(24785)

H0: mean = 1500 Claim
H1: mean > 1500
Sample mean 1509.5
Std error of mean 4.854
t-test Using 95% CI, (right handed only: .05)
t = 9.5/(4.854/sqrt(25)) = 9.5/.9708
P(t > 9.78574)= ~ 0, basically 'off the charts'
reject H0

Answer by ed42ptt(4)

(X ̅-μ)/(σ/√n)=>Standard Error=σ/√n=(24.27)/√25=4.854;
Thus using (X ̅-μ)/E would be the new formula to use (±1509.5-1500)/(4.854)=>(±9.5)/(4.854)≈ (-1.957 and 1.957) are the Smaller and larger Critical Values. In addition, just in case you need the p-value for 2 tails it would be 0.062 and for one tail it would be 0.031. You can used the Free p-value online calculator to calculate these additional values and double check that they are correct.